# ReQISC: A Reconfigurable Quantum Computer Microarchitecture and Compiler Co-Design

### v1

@@ 4.2. Addressing the “near-identity” issue
Following  Algorithm 1, we find that realizing gates with Weyl coordinate close to the identity gate (common in QFT and Hamiltonian simulation kernels) in optimal time requires unbounded pulse amplitudes, which are experimentally infeasible. While trading gate time for bounded amplitudes is possible (chen2024one), it leads to a more error-prone implementation due to longer duration and complex control. To address this, we propose transforming the near-identity gate into its mirror symmetry (cross2019validating) differing by a $\mathrm{SWAP}$ gate, as
$\mathrm{SWAP}\cdot\mathrm{Can}(x,y,z)\sim\begin{cases}\mathrm{Can}\left(\frac{\pi}{4}-z,\frac{\pi}{4}-y,x-\frac{\pi}{4}\right),&\textrm{if }z\geq 0\\ \mathrm{Can}\left(\frac{\pi}{4}+z,\frac{\pi}{4}-y,\frac{\pi}{4}-x\right),&\textrm{if }z<0\end{cases}.$
Thus the mirrors of near-identity gates are located far away from the identity, which is illustrated in Figure 4 (a). In the compilation stage, we appended the $\mathrm{SWAP}$ to each near-identity 2Q gate, leaving the circuit rewiring effects of these $\mathrm{SWAP}$s to be resolved at compile time. Specifically, we define a coordinate norm threshold to determine if a 2Q gate is near-identity and requires mirroring. The threshold $r$ in general depends on hardware parameters (chen2024one). The impacts of inserted $\mathrm{SWAP}$s are tracked by recording the altered qubit mappings, which are then used to rewire the entire circuit at the final step. Thus we resolve the issue at compile time without introducing additional #2Q overhead.
@@ 4.3. Hardware implementation
To implement our presented microarchitecture on any quantum platforms, the only requirement is that $H_{1},H_{2}$ can be applied simultaneously with the 2Q interaction. In the case of flux-tunable superconducting qubits, requirements for the control equipment are the same as that of typical experiments, i.e., standard microwave drives, AWGs, and mixers are sufficient, as recently validated experimentally (chen2025efficient). Given these minimal requirements, many platforms, e.g., flux-tunable transmons (arute2019quantum; acharya2024quantum), fluxonium qubits (nguyen2019high; bao2022fluxonium), Floquet qubits (nguyen2024programmable), semiconductor spin qubits (dijkema2025cavity), trapped ions (yale2025realization), and some neutral-atom devices (barredo2015coherent) could immediately benefit from this scheme. Herein we highlight its broad applicability via case studies and give the detailed pulse-control description.
[…]
(3) $H=-\frac{\omega_{1}}{2}ZI-\frac{\omega_{2}}{2}IZ+gYY+V_{1}(t)YI+V_{2}(t)IY,$
where $\omega_{i}$ are the frequencies of the qubits, $g$ is the coupling strength, and $V_{i}(t)$ are the local drives. By conjugation by a local unitary, we can transform this to
@@
which falls under our framework. We can simply tune different qubit frequencies to control the $ZI,IZ$ terms while local drives can be square pulses with variable amplitudes instead of sinusoidal pulses to control the $XI,IX$ terms. This Hamiltonian example is explicitly not covered by (chen2024one) but can appear in hardware devices. Our gate scheme can handle this along with any other coupling Hamiltonian.
Through Algorithm 1, we can derive the subschemes (ND, EA+, or EA-) and control parameters for different situations, as shown in Figure 4 (c-e). By means of our scheme, some takeaways for ISA design are worth noting. For example, gates requiring no or weak local drives (low amplitude) potentially achieve higher fidelities than those requiring stronger drives (e.g., $\mathrm{CNOT}$ under $\mathrm{XX}$ coupling v.s. $\mathrm{CNOT}$ under $\mathrm{XY}$ coupling, $\mathrm{iSWAP}$ under $\mathrm{XY}$ coupling v.s. $\mathrm{iSWAP}$ under $\mathrm{XX}$ coupling), since precise calibration of local drives with the coupling turned on is nontrivial in practice. Using our gate scheme, $\mathrm{SWAP}$ is far less expensive than the conventional three-$\mathrm{CNOT}$ decomposition, as shown for $\mathrm{XY}$ coupling in Figure 4 (b). Even the $\mathrm{CNOT}$ implementation via our scheme is 1.411.41x faster than the traditional control scheme (krantz2019quantum), that is, $\pi/2g$ v.s. $\pi/\sqrt{2}g$. This speedup is critical for suppressing errors on platforms where decoherence is the dominant error source. Besides, our scheme does not introduce the additional $\lvert 11\rangle\leftrightarrow\lvert 02\rangle$ transition, thereby minimizing leakage error and dynamic stray $\mathrm{ZZ}$ crosstalk. Part of these features have been validated in superconducting transmon qubits by chen2025efficient, where versatile 2Q gates are implemented in high fidelity (on average 99.3799.37%), with optimal gate duration and simple pulse control. As our scheme represents a generalized approach to directly implement arbitrary $\mathrm{SU}(4)$ on physical platforms with other coupling Hamiltonians, it poses more opportunities for hardware-native ISA design and compiler optimization.
@@ 4.4. Gate calibration
Gate calibration can follow a similar procedure as outlined in (chen2025efficient), where $\mathrm{XY}$ coupling is assumed. First, the $\mathrm{iSWAP}$-family component of the Hamiltonian (coupling term $H$) and the drive components ($H_{1}$ and $H_{2}$) are separately calibrated to determine how the model parameters $g$, $\Omega_{1,\,2}$, and $\delta$ depend on physical control parameters. This provides an initial estimate for physical control parameters. Then, the two parts of the Hamiltonian are simultaneously applied. Further optimization is achieved by first applying quantum process tomography (greenbaum2015introduction). This is used to measure the Weyl coordinate of the physical gate realized. Control parameters are tuned to minimize the Euclidean distance from target coordinates. Further fine tuning of the parameters can be achieved via cross-entropy benchmarking (arute2019quantum).
Such a calibration method is readily generalizable. For example, instead of calibrating an $\mathrm{iSWAP}$-family gate, one would calibrate an $\mathrm{XX}(\theta)$ rotation for the case of $\mathrm{XX}$ coupling. To calibrate a continuous gate set, the idea in (chen2024one) could be utilized. In this case, we can calibrate the mapping between the control parameters and model parameters. Such a mapping can be expressed as some parameterized mathematical function. We can then optimize the parameters of this function via fully randomized benchmarking (Haar-random 2Q gates) (kong2021framework) or other more general benchmarking schemes (chen2022randomized) with respect to different gate distributions.
@@ 5.1. Local optimization: Hierarchical synthesis
We first attempt to exploit $\mathrm{SU}(4)$’s synthesis potential based on its resource-bound analysis and work out a hierarchical synthesis structure to perform aggressive local optimization at the circuit level. We introduce a “compactness” metric across local subcircuits to guide co-optimization and devise a supporting DAG compacting pass with the approximate commutation rules in our findings.
@@ 5.1.1. Resource analysis and approximate synthesis with $\mathbf{SU}(4)$
The lower bound of 2Q basis gate count given a quantum ISA typically reflects its intrinsic synthesis power. The bound for an $\mathrm{SU}(4)$-based ISA is $b_{\mathrm{SU}(4)}(n)=\lceil(4^{n}-3n-1)/9\rceil$, implying a theoretical 55.6% reduction in #2Q gates compared to a $\mathrm{CNOT}$-based ISA, whose bound is $b_{\mathrm{CNOT}}(n)=\lceil(4^{n}-3n-1)/4\rceil$. While no constructive algorithms can achieve this bound, the approximate synthesis technique has demonstrated this capability (khatri2019quantum; madden2022best). Through structural optimization of gate arrangements and numerical optimization of gate parameters, approximate synthesis can generate circuits with fewer gates that are equivalent to a target unitary within a permissible precision (davis2019heuristics). This precision is typically set to be negligible ($^{-15}-1010$$^{-10}$, measured by infidelity $1-\frac{1}{N}\lvert\mathrm{Tr}(U^{\dagger}V)\rvert$) compared to physical errors, making the synthesis effectively exact for practical purposes.
(a) Example and pipeline of hierarchical synthesis (HS). (a) The original alu-v2_33 circuit is partitioned into 3 three-qubit blocks with #2Q equal to 1, 8, and 8. After approximate synthesis on the latter two blocks, each has 5 $\mathrm{SU}(4)$s. The overall #2Q is reduced from 17 to $1$. (b) HS involves two-tier partitioning, followed by DAG compacting and conditional approximate synthesis.
Figure 6. Exchanging two “approximately commutative” $\mathrm{SU}(4)$’s improves compactness. This pattern occurs in grover_5_0. Different colors indicate that the parameters of $\mathrm{SU}(4)$ or $\mathrm{SU}(8)$ have changed.
@@ 5.1.2. Hierarchical synthesis
To leverage the theoretical synthesis potential of $\mathrm{SU}(4)$ and apply approximate synthesis to large-scale real-world programs, we employ a two-tier hierarchical synthesis approach. It first partitions the original circuit into 2Q blocks, each of which is fused into an $\mathrm{SU}(4)$ gate. Then, it partitions the circuit, now consisting solely of $\mathrm{SU}(4)$ gates, into $w$-qubit blocks to determine if each block can be synthesized with fewer $\mathrm{SU}(4)$ gates through approximate synthesis. Specifically, our compiler conditionally performs approximate synthesis blocks whose $\mathrm{SU}(4)$ number is greater than a threshold $m_{th}$. Thus, two hyperparameters need to be carefully considered to strike a trade-off between computational overhead and 2Q gate count reduction effect via approximate synthesis: partitioning granularity $w$ and 2Q gate count threshold $m_{th}$. We determine optimal default values through empirical analysis: (1) Partition granularity ($w=3$): Approximate synthesis is computationally expensive, scaling exponentially with circuit width. We found that 3Q partitioning offers a sweet spot, as 3Q subcircuits frequently contain more gates than the theoretical minimum ($b_{\mathrm{SU}(4)}(3)=6$), providing ample optimization opportunities. In contrast, 4Q partitioning results in too slow computation of underlying approximate synthesis, and its potential for reduction is less frequently realized in practice, as $b_{\mathrm{SU}(4)}(4)=27$ is too large. (2) Synthesis threshold ($m_{th}=4$): In principle, $m_{th}$ can be set to the #2Q lower bound, i.e., $6$. However, empirical observations indicate that as few as $5$ or even $4$ gates are often sufficient to approximate 3Q blocks within the permissible precision, suggesting that a more aggressive $m_{th}$ can yield greater gate count reductions. As shown in Figure 6(a), this method reduces the #2Q for the alu-v2_33 circuit from $17$ to $11$.
@@ 5.1.3. Compactness and DAG compacting
Under the guidance of the compactness metric and the gate scheduling idea, we invent the DAG compacting pass to conduct aggressive optimization for partitioned $\mathrm{SU}(4)$-based subcircuits that comprise a nested DAG. The DAG compacting pass exploits opportunities of exchanging “commutative” $\mathrm{SU}(4)$s to enhance compactness. For example, a gate sequence like $\mathrm{SU}(4)_{1,2},\,\mathrm{SU}(4)_{2,3}]$ can often be exchanged into $\mathrm{SU}(4)^{\prime}_{2,3},\,\mathrm{SU}(4)^{\prime}_{1,2}]$ with negligible error (Figure 6). By performing this commutation, $\mathrm{SU}(4)^{\prime}_{1,2}$ can merge with the subsequent 3Q block $\mathrm{SU}(8)_{0,1,2}$, thereby increasing compactness. The pass iteratively applies these feasible exchanges to maximize the compactness-increasing opportunities, enabling more aggressive optimization.
@@ 5.2. Global optimization: Program-aware synthesis
The hierarchical synthesis works at the circuit level and will induce diverse parameterized quantum gates thus with uncontrollable calibration overhead. It is impractical for large-scale program compilation on its own. Thus, an approach to globally optimize quantum programs and managing the calibration overhead is necessary.
@@ 5.2.1. Real-world program patterns
In general, real-world quantum algorithms can be categorized into two types: (1) those solving classical problems, that is, quantum versions of digital logic functions, constructed through binary to qubit encoding, and (2) those solving more quantum or optimization problems, typically constructed through Hamiltonian simulation. For the latter type of programs, there are already highly-effective global compilation techniques that can be seamlessly applied to the $\mathrm{SU}(4)$ ISA, especially for those ISA-independent dedicated compilers such as Rustiq (debrugière2024faster), QuCLEAR (liu2025quclear), and Phoenix (yang2025phoenix). Therefore, we focus more on how to globally optimize the first type of quantum programs.
Specifically, these programs are composed of $\mathrm{CX}$$\mathrm{CCX}$$\mathrm{MCX}$ gates (sometimes with 1Q rotations and Paulis), as exemplified in Figure 7. This feature is ideal for a template-based synthesis strategy—first find the optimal synthesis schemes for these high-level semantics and then unroll them within circuits. We refine 3Q blocks as the IRs and constructs their optimal synthesis templates again via approximate synthesis, since intricate components such as Peres (thapliyal2009design) and MAJ/UMA (cuccaro2004new) are usually represented as 3Q circuits. Moreover, 3Q represents the perfect granularity for efficient yet powerful synthesis as discussed in Section 5.1.2. For $\mathrm{MCX}$, it can be first recursively decomposed into $\mathrm{CCX}$ gates (barenco1995elementary). This approach simultaneously enables effective optimization and benefits calibration efficiency by reusing a finite set of highly optimized building blocks.
@@
Figure 8. Template-based synthesis building on a pre-synthesized template library and the selective assembly strategy.

### v2

@@ 4.3. Addressing the “near-identity” issue
Following  Algorithm 1, realizing gates with Weyl coordinates close to the origin (common in quantum Fourier transform (QFT) and Hamiltonian simulation kernels) in optimal time requires unbounded pulse amplitudes, which are experimentally infeasible. While trading gate time for bounded amplitudes is possible (Chen et al., 2024), the longer gate time is still undesirable. To address this, we propose transforming the near-identity gate into its mirror gate (Cross et al., 2019) which differs by a $\mathrm{SWAP}$ gate:
$\mathrm{SWAP}\cdot\mathrm{Can}(x,y,z)\sim\begin{cases}\mathrm{Can}\left(\frac{\pi}{4}-z,\frac{\pi}{4}-y,x-\frac{\pi}{4}\right),&\textrm{if }z\geq 0\\ \mathrm{Can}\left(\frac{\pi}{4}+z,\frac{\pi}{4}-y,\frac{\pi}{4}-x\right),&\textrm{if }z<0\end{cases},$
and subsequently altering qubit mapping at compile time. It is clear that the Weyl coordinates of the mirrors of near-identity gates are far from the origin, as depicted in Figure 5(a). During compilation, we append a $\mathrm{SWAP}$ to each near-identity 2Q gate, leaving the circuit rewiring effects of these $\mathrm{SWAP}$s to be resolved at compile time. Specifically, we define a coordinate norm threshold to determine if a 2Q gate is near-identity and requires mirroring. The threshold $r$ in general depends on hardware parameters (Chen et al., 2024). The impacts of inserted $\mathrm{SWAP}$s are tracked by recording the altered qubit mappings. Thus we resolve the issue at compile time without necessarily introducing additional 2Q gate count overhead. For example, Figure 5(c) illustrates an example for the four-qubit QFT circuit, which eventually leads to a circuit with no extra 2Q gates and only an updated final mapping to track.
@@ 4.4. Hardware implementation
Figure 6. Hardware implementation of the ReQISC microarchitecture. (a) Gate time landscape under $\mathrm{XY}$ coupling ($H=\frac{g}{2}(XX+YY)$); circle size indicates gate duration. (b-c): Pulse control subschemes, under $\mathrm{XY}$ and $\mathrm{XX}$ couplings, respectively. (d) Local drive amplitudes required for representative gate families. The scaling factor $s$ is used to take fractions of gates. For example, $\mathrm{B}^{s}\sim\mathrm{Can}(\frac{\pi}{4}s,\frac{\pi}{8}s,0)$. The $\mathrm{iSWAP}$ family requires no local drives so it is not shown in the figure; the $\mathrm{CNOT}$ family and $\mathrm{B}$ family require only one-side drive ($A_{1}\neq 0$); $\mathrm{SWAP}$ family requires both-side drives ($A_{1}=A_{2}\neq 0$). Frequency-related quantities ($A_{1}$, $A_{2}$, $\delta$) are normalized by the coupling strength $g$.
To implement our presented microarchitecture on a hardware platform, the only requirement is that $H,H_{1},H_{2}$ can be applied simultaneously. This is true for the case of flux-tunable superconducting qubits: standard microwave drives, arbitrary wave generators, and mixers are sufficient, as recently shown in an experiment (Chen et al., 2025). Given these minimal requirements, many platforms, e.g., flux-tunable transmons (Arute et al., 2019; Acharya et al., 2024), fluxonium qubits (Nguyen et al., 2019; Bao et al., 2022), Floquet qubits (Nguyen et al., 2024), semiconductor spin qubits (Dijkema et al., 2025), trapped ions (Yale et al., 2025), and some neutral-atom devices (Barredo et al., 2015) could immediately benefit from this scheme.
We outline how to realize common 2Q gates ($\mathrm{CNOT}$, $\mathrm{iSWAP}$, $\mathrm{B}$, etc.) with two typical coupling Hamiltonians: $\mathrm{XY}$ coupling ($H=\frac{g}{2}\,(XX+YY)$) and $\mathrm{XX}$ coupling (H=gXXH=g\,$X$). For instance, the former is common in flux-tunable superconducting qubits (Krantz et al., 2019; Arute et al., 2019); the latter is the dominant coupling type for trapped ions (Yale et al., 2025). The latter type of coupling can also appear in capacitively coupled flux-tunable transmons by working in the lab frame instead of the drive frame (Krantz et al., 2019):
@@
where $\omega_{i}$ are the frequencies of the qubits, $g$ is the coupling strength, and $V_{i}(t)$ are the local drives. We can simply tune different qubit frequencies to control the $ZI,IZ$ terms while local drives can be square pulses with variable amplitudes instead of sinusoidal pulses to control the $XI,IX$ terms. This Hamiltonian is not supported by previous work (Chen et al., 2024), yet can arise for certain hardware devices. Our gate scheme can handle this or any other coupling Hamiltonian.
Through Algorithm 1, we can derive the subschemes (ND, EA+, or EA-) and control parameters as shown in Figure 6(b-d). With respect to our gate scheme, there are some takeaways for ISA design. For example, gates requiring no or low amplitude drives can potentially achieve higher fidelities than those requiring stronger drives (e.g., $\mathrm{CNOT}$ under $\mathrm{XX}$ coupling vs. $\mathrm{CNOT}$ under $\mathrm{XY}$ coupling, $\mathrm{iSWAP}$ under $\mathrm{XY}$ coupling vs. $\mathrm{iSWAP}$ under $\mathrm{XX}$ coupling, respectively). This is because precise calibration of local drives with the coupling turned on is nontrivial in practice. Using our gate scheme, $\mathrm{SWAP}$ is far less expensive than the conventional three-$\mathrm{CNOT}$ decomposition, as shown for $\mathrm{XY}$ coupling in Figure 6(b). Even the $\mathrm{CNOT}$ implementation via our scheme is 1.411.41x faster than the traditional control scheme on superconducting processors (Krantz et al., 2019), that is, $\pi/2g$ vs. $\pi/\sqrt{2}g$. This speedup is critical for suppressing errors on platforms where decoherence is the dominant error source. Besides, our scheme does not introduce the additional $\lvert 11\rangle\leftrightarrow\lvert 02\rangle$ transition, thereby minimizing leakage error and dynamic stray $\mathrm{ZZ}$ crosstalk. These advantages were demonstrated on transmons in (Chen et al., 2025), where various 2Q gates are implemented in high fidelity (on average 99.3799.37%), with optimal gate duration. As our scheme represents a general approach to directly implement arbitrary $\mathrm{SU}(4)$ on physical platforms with arbitrary coupling Hamiltonians, it provides more opportunities for hardware-native ISA design and compiler optimization.
@@ 4.5. Gate calibration
Gate calibration can follow a similar procedure as outlined in (Chen et al., 2025), where $\mathrm{XY}$ coupling is assumed. First, the $\mathrm{iSWAP}$-family component of the Hamiltonian (coupling term $H$) and the drive components ($H_{1}$ and $H_{2}$) are separately calibrated to determine how the model parameters $g$, $\Omega_{1,\,2}$, and $\delta$ depend on physical control parameters. This provides an initial estimate for physical control parameters. Then, the two parts of the Hamiltonian are simultaneously applied. Further optimization is achieved by first applying quantum process tomography (Greenbaum, 2015). This is used to measure the Weyl coordinate of the physical gate realized. Control parameters are tuned to minimize the Euclidean distance from target coordinates. Further fine-tuning of the parameters can be achieved via cross-entropy benchmarking (Arute et al., 2019).
Such a calibration method is readily generalizable. For example, instead of calibrating an $\mathrm{iSWAP}$-family gate, one would calibrate an $\mathrm{XX}(\theta)$ rotation for the case of $\mathrm{XX}$ coupling. To calibrate a continuous gate set, the idea in (Chen et al., 2024) could be utilized. In this case, we can calibrate the mapping between the control parameters and model parameters. Such a mapping can be expressed as some parameterized mathematical function. We can then optimize the parameters of this function via fully randomized benchmarking (Haar-random 2Q gates) (Kong, 2021) or other more general benchmarking schemes (Chen et al., 2022) with respect to different gate distributions.
@@ 5.1. Local optimization: Hierarchical synthesis
We first attempt to exploit $\mathrm{SU}(4)$’s synthesis potential based on its resource-bound analysis and devise a hierarchical synthesis structure to perform aggressive local optimization at the circuit level. We introduce a “compactness” metric across local subcircuits to guide co-optimization and devise a supporting DAG compacting pass with the approximate commutation rules in our findings.
@@ 5.1.1. Resource analysis and approximate synthesis with $\mathbf{SU}(4)$
The lower bound of 2Q basis gate count given a quantum ISA typically reflects its intrinsic synthesis power. The bound for an $\mathrm{SU}(4)$-based ISA is $b_{\mathrm{SU}(4)}(n)=\lceil(4^{n}-3n-1)/9\rceil$, implying a theoretical 55.6% reduction in #2Q gates compared to a $\mathrm{CNOT}$-based ISA, whose bound is $b_{\mathrm{CNOT}}(n)=\lceil(4^{n}-3n-1)/4\rceil$. While no constructive algorithms can achieve this bound, the approximate synthesis technique has demonstrated this capability (Khatri et al., 2019; Madden and Simonetto, 2022). Through structural optimization of gate arrangements and numerical optimization of gate parameters, approximate synthesis can generate circuits with fewer gates that are equivalent to a target unitary within a permissible precision (Davis et al., 2019). This precision is typically set to be negligible ($^{-15}-1010$$^{-10}$, measured by infidelity $1-\frac{1}{N}\lvert\mathrm{Tr}(U^{\dagger}V)\rvert$) compared to physical errors, making the synthesis effectively exact for practical purposes.
(a) HS example for alu-v2_33 circuit
(b) HS pipeline
Figure 7. Example and pipeline of hierarchical synthesis (HS). (a) The original alu-v2_33 circuit is partitioned into 3 three-qubit blocks with #2Q equal to 1, 8, and 8. After approximate synthesis on the latter two blocks, each has 5 $\mathrm{SU}(4)$s. The overall #2Q is reduced from 17 to $1$. (b) HS involves two-tier partitioning, followed by DAG compacting and conditional approximate synthesis.
Figure 8. Exchanging “approximately commutative” $\mathrm{SU}(4)$s improves compactness. This pattern occurs in grover_5_0. Different colors indicate that parameters of $\mathrm{SU}(4)$ or $\mathrm{SU}(8)$ have changed.
@@ 5.1.2. Hierarchical synthesis
To leverage the theoretical synthesis potential of $\mathrm{SU}(4)$ and apply approximate synthesis to large-scale real-world programs, we employ a two-tier hierarchical synthesis approach. It first partitions the original circuit into 2Q blocks, each of which is fused into an $\mathrm{SU}(4)$ gate. Then, it partitions the circuit, now consisting solely of $\mathrm{SU}(4)$ gates, into $w$-qubit blocks to determine if each block can be synthesized with fewer $\mathrm{SU}(4)$ gates through approximate synthesis. Specifically, our compiler conditionally performs approximate synthesis on blocks whose $\mathrm{SU}(4)$ count is greater than a threshold $m_{th}$. Thus, two hyperparameters need to be carefully considered to strike a trade-off between computational overhead and 2Q gate count reduction effect via approximate synthesis: partitioning granularity $w$ and 2Q gate count threshold $m_{th}$. We determine optimal default values through empirical analysis: ① Partition granularity ($w=3$): Approximate synthesis is computationally expensive, scaling exponentially with circuit width. We found that 3Q partitioning offers an optimal balance, as 3Q subcircuits frequently contain more gates than the theoretical minimum ($b_{\mathrm{SU}(4)}(3)=6$), providing ample optimization opportunities. In contrast, 4Q partitioning results in excessive computational overhead for the underlying approximate synthesis, and its potential for reduction is less frequently realized in practice, as $b_{\mathrm{SU}(4)}(4)=27$ is too large. ② Synthesis threshold ($m_{th}=4$): In principle, $m_{th}$ can be set to the #2Q lower bound, i.e., $6$. However, empirical observations indicate that as few as $5$ or even $4$ gates are often sufficient to approximate 3Q blocks within the permissible precision, suggesting that a more aggressive $m_{th}$ can yield greater gate count reductions. As shown in Figure 7, this method reduces the #2Q for the alu-v2_33 circuit from $17$ to $11$.
@@ 5.1.3. Compactness and DAG compacting
Under the guidance of the compactness metric and the gate scheduling idea, we propose the DAG compacting pass to conduct aggressive optimization for partitioned $\mathrm{SU}(4)$-based subcircuits that comprise a nested DAG. The DAG compacting pass exploits opportunities of exchanging “commutative” $\mathrm{SU}(4)$s to enhance compactness. For example, a gate sequence like $\mathrm{SU}(4)_{1,2},\,\mathrm{SU}(4)_{2,3}]$ can often be exchanged into $\mathrm{SU}(4)^{\prime}_{2,3},\,\mathrm{SU}(4)^{\prime}_{1,2}]$ with negligible error (Figure 8). By performing this commutation, $\mathrm{SU}(4)^{\prime}_{1,2}$ can merge with the subsequent 3Q block $\mathrm{SU}(8)_{0,1,2}$, thereby increasing compactness. The pass iteratively applies these feasible exchanges to maximize the compactness-increasing opportunities, enabling more aggressive optimization.
@@ 5.2. Global optimization: Program-aware synthesis
The hierarchical synthesis works at the circuit level and will induce diverse parameterized quantum gates, leading to uncontrollable calibration overhead. It is impractical for large-scale program compilation on its own. Thus, an approach to globally optimizing quantum programs and managing the calibration overhead is necessary.
@@ 5.2.1. Real-world program patterns
In general, real-world quantum algorithms can be categorized into two types: (1) those solving classical problems, i.e., quantum versions of digital logics, constructed through binary to qubit encoding, and (2) those solving more quantum or optimization problems, typically constructed through Hamiltonian simulation. For the latter type of programs, there are already highly-effective global compilation techniques that can be seamlessly applied to the $\mathrm{SU}(4)$ ISA, especially for those ISA-independent dedicated compilers such as Rustiq (debrugière2024faster), QuCLEAR (Liu et al., 2025), and PHOENIX (Yang et al., 2025). Therefore, we focus on how to globally optimize the first type of quantum programs.
Specifically, these programs are composed of $\mathrm{CX}$$\mathrm{CCX}$$\mathrm{MCX}$ gates (sometimes with 1Q rotations and Paulis), as exemplified in Figure 9. This feature is ideal for a template-based synthesis strategy—first find the optimal synthesis schemes for these high-level semantics and then unroll them within circuits. We refine 3Q blocks as the IRs and construct their optimal synthesis templates again via approximate synthesis, since intricate components such as the Peres gate (Thapliyal and Ranganathan, 2009) and MAJ/UMA (Cuccaro et al., 2004) are usually represented as 3Q circuits. Moreover, 3Q represents the perfect granularity for efficient yet powerful synthesis as discussed in Section 5.1.2. For $\mathrm{MCX}$, it can be first recursively decomposed into $\mathrm{CCX}$ gates (Barenco et al., 1995). This approach simultaneously enables effective optimization and benefits calibration efficiency by reusing a finite set of highly optimized building blocks.
@@
Figure 10. Template-based synthesis building on a pre-synthesized template library and selective assembly.